EDITORIAL

Terry Goodman

Welcome to the first issue of the Missouri Journal of Mathematical Sciences. Those of us who serve as the editorial board are excited about the journal and the possibilities for it in the future. We want MJMS to provide an opportunity for those in the mathematical sciences to share ideas and information. While MJMS is but one among many journals for the mathematical sciences, we believe that there is always a need for an exchange of ideas and that this journal will help to meet that need, especially in Missouri.

We intend to make MJMS of interest to a wide range of readers. This first issue illustrates that variety. The articles by Russell Euler and Shing-Seung So focus on some interesting results in special functions and topological spaces, respectively. Stan Hartzler expresses some strong observations with a point of view that will generate thought and discussion, both pro and con. Sherralyn Craven shares with us some information and insights about the sixth International Congress on Mathematical Education. Curtis Cooper has provided the reader with some very interesting and challenging problems. We look forward to seeing your solutions for them. Finally, we have included a listing of professional meetings and seminars for the current academic year. While we certainly will not try to provide “something for everyone”, we do hope that mathematics teachers at many levels will find MJMS to be useful and interesting.

We encourage you to participate in MJMS. Certainly, we hope you will order a subscription to the journal. An order form for individual and institutional subscriptions is found on page 31. We invite you to submit articles, commentaries, problems and announcements. Guidelines for contributors are found in the inside back cover. You may be interested in serving as a referee for MJMS. If so, you should contact Robert Kennedy or William Grimes at the enclosed address. We also would welcome letters concerning your ideas and suggestions for MJMS as well as your response to any specific article. Finally, we hope that you will share this issue with your colleagues and encourage them to participate.

Producing a journal is a major undertaking. Before this issue—we did not realize how major! We appreciate the help and support of Edward Davenport, Chair, and other members of the Department of Mathematics and Computer Science at Central Missouri State University.

The financial assistance provided by the University Professional Enhancement Committee and by Benjamin Taggie, Dean of the College of Arts and Sciences has made this journal possible. We thank them for their support.

We also thank Ellen Mowatt for her design of the logo for MJMS and Teresa Goodman for her assistance with the artwork for the cover.

The input from various referees was also extremely helpful and we thank them for their contributions.

A special note of thanks needs to go to Becky White for her help in typing and organizing the journal. We appreciate her ability to decipher our writing, meet our deadlines, learn new ideas, and correct our mistakes.